A Complete Bibliography of Publications in ACM SIGSOFT Software Engineering Notes: 2100–2109

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

04 April 2024 Version 1.03

Title word cross-reference

4 [NSBC^{+}23].

1 [Sch22a]. 11th [SABB24]. 12th [SPC22]. 13th [KMS23]. ’19 [VFHD20, BFS22, FM21, Kam23a, KS23, MBL21]. 1st [SPC22, DP23, JTNBC24, TRH^{+}20, VFHD20].

2 [Sch22b]. ’20 [VHFD21]. 2019 [APA^{+}20, Niu20]. 2020 [LWP^{+}20, Pfa21b, Pfa21a, Pfa21c]. 2021 [Bro21]. 2022 [SNA^{+}23, Sol22a, Sol22b]. 2023 [AAC^{+}24, LNP^{+}23, LCA^{+}23, SABB24, Sol23a, Sol23b, Sol23c]. ’21 [HPSV22]. ’22 [HPSV23]. 2nd [CLG^{+}22, CSS20, PD23, RDF^{+}20, SWG21, VHFD21].

3 [Sch23b]. 38th [AAC^{+}24]. 3rd

[ECPJ^{+}23, HPSV22].

4 [Sch23a]. 4th [ESGJ^{+}23, HPSV23].

5th [AAC^{+}24, WHW23].

6th [KFG^{+}21].

7th [XKD22].

8th [GMD24, YAT^{+}21].

A-TEST [KMS23, VPE^{+}20]. Academia [MBL21]. Academic [Sch22a, Sch22b, Sch23b, Sch23a]. ACM [AAC^{+}24, Pfa21b, Pfa21a, Pfa21c, RR20, Ral21, SGHJ20, WHW23, SABB24].

ACM/IEEE [SABB24]. Action [Mau23]. Adaptation [WGB^{+}22]. Adaptive [WCM^{+}23]. Adoption [ED22a]. Advanced [HPSV22, HPSV23, VFHD20, VHFD21]. Advancing [BCL^{+}23b]. Affects [ED23a]. Africa [BHC20, BHC21]. After [BFS22, ED23a, MBL21]. Age [PPPC20]. Agenda [WCM^{+}23]. Agile [XdSBT24]. AI [ED22a, ED23c, ED24, NSJ^{+}22, NSBC^{+}23]. American [SPC22]. Announcing [Di 20]. Approaches [dS23]. Architecture [Ham23]. Architectures [LCA^{+}23, RDF^{+}20]. Artifacts [WGB^{+}22]. Artificial [HPSV22, HPSV23, Sch20, VFHD20, VHFD21, YAT^{+}21]. ASE [AAC^{+}24, Niu20]. ASE4Games [CPGP22]. Assessment [Mba23]. ASYDE [AAC^{+}24]. Attitudinal [Kam23a]. Australasia [LTG^{+}21]. Automated [AAC^{+}24]. Automating [KMS23, VPE^{+}20]. Autonomous [RDF^{+}20]. Award [Shi21]. Awardee [Sol22a, Sol22b, Sol23a, Sol23b, Sol23c]. Awardees [Niu20]. aware [RDF^{+}20]. Awareness [FU21].

Based [BBC^{+}20, DP23, PD23]. be [PGM20]. Because [Sch23c]. Behind [ED22c, ED24]. better [PGM20]. biblatex [Di 20]. biblatex-software [Di 20]. Binary [HKVD21]. Blended [LCA^{+}23, Moh23]. blind [PGM20]. Blog [PM21]. Bobby [Rog22a]. Book [ECN20, Rog22a, Rog22b]. Bots [SWG21, WGW22]. BotSE [SWG21, WGW22]. Brazilian [SPC22]. Briand [Pfa21b]. Bridging [MZHS21]. BRIGHT [BHC20]. bug [PGG21, RK21]. Building [BHC20, Rog22a]. Business [SGHJ20, WHW23, WHW23].

Can [ED23b, ED23c]. Capacity [BHC20]. CAPS [Niu20]. Case [KMS23, VPE^{+}20]. CE [Leh23]. Centric [CLG^{+}22, Med21]. CEO [ED24]. Challenges [ED22d, FTP^{+}23]. Changes [Kam23a, PGG21]. Chaos [FM21]. Checking [PK23, WKJ^{+}23]. Chen [Sol23a]. Chief [HX20]. China [Sch23a]. Chunyang [Sol23a]. churn [PGG21]. citation [Di 20]. Clone [HKVD21]. Cloud [ECN20]. co [AAC^{+}24, KMS23, HX20]. Co-Editors-in-Chief [HX20]. co-located [AAC^{+}24, KMS23]. Code [HKVD21, WKJ^{+}23, PGG21]. Collaboration [BCL^{+}23b]. Collaborations [FM22, MBL21]. Combinatorial [Tzo20]. Committee [TRK^{+}21]. community [PGM20]. Companies [ED23d]. Company [FM22]. comparative [RK21]. Comparing [PGG21]. Compass [ZA21]. Compliance [Sch22b]. Component [Hei23]. composite [RK21]. Comprehension [Hei23, Tzo20]. Computer [Ven22]. Computing [BBC^{+}20, ECN20, FTP^{+}23, PPPC20]. Conference [SPC22, KMS23]. Configuration [CS20]. Context [Hei23, RDF^{+}20]. Context-aware [RDF^{+}20]. Contexts [BCL^{+}23b]. Continuous [KFG^{+}21, Leh23]. COVID [BFS22, FM21, Kam23a, KS23, MBL21]. COVID-19 [BFS22, FM21, Kam23a, KS23, MBL21]. crazy [Sch23c]. Create [ED22b]. Critical [ECPJ^{+}23, ESGJ^{+}23]. cross [Mau23]. cross-platform [Mau23]. Cubesats [Sch23b]. Currently [ED22d]. Customer [ED22b]. Cyber [CLG^{+}22, NSJ^{+}22, NSBC^{+}23]. Cyber-Physical [NSJ^{+}22, NSBC^{+}23]. Cybersecurity [ECPJ^{+}23, ESGJ^{+}23].

Data [Jah22, NSJ^{+}22, NSBC^{+}23]. Datafication [Jah22]. Debt [XdSBT24]. Decade [SCAB23]. Deemed [Sch22a, Sch22b, Sch23b]. Deep [BCS^{+}23]. DeepTest [BCS^{+}23]. Dependencies [SBT^{+}20]. Deploying [Rog22a].

Deployment [MGW20]. Design [ECN20, KMS23, MGW20, VPE^{+}20, Rog22b]. Designing [Rog22a]. Detection [HKVD21]. Developers [PM21]. DEvelopment [AAC^{+}24, BCL^{+}23b, Moh23]. Differences [GP23]. Different [Kam23b]. digital [Mau23]. Dilemma [ED22c, ED24]. Dimensions [FM23]. Dismissal [ED24]. Dissertation [Shi21]. Distant [FM22]. Diversity [FM23, JJL23]. Do [ED23c, PM21, Sch20]. Doctoral [Shi21]. Does [ED23c]. Domain [Rog22b]. Domain-Driven [Rog22b]. Double [PGM20]. Double-blind [PGM20]. DQ’22 [NSBC^{+}23]. Driven [Rog22b, AAY^{+}21b, San23]. Drivers [ED22c]. During [MBL21].

EASEAI [HPSV22, HPSV23, VFHD20, VHFD21]. East [BHC21]. easy [Di 20]. Ecosystems [SCAB23, SABB24, SGHJ20]. Editions [VPE^{+}20, SBT^{+}20]. Editors [HX20]. Education [HPSV22, HPSV23, Kam23b, KT23, KBT23, PK22, TFTS20, VFHD20, VHFD21, Ven22]. Educator [Wil20]. Effects [Mba23]. Efficiency [Shi21]. Efficient [PK23]. Electric [Sch23c]. Elements [ED22b]. Emergency [DMM21]. Emerging [BHC21, TFTS20]. EMIP [ABS23]. Emotion [FU21]. EmpiRE [GMD24]. Empirical [GMD24, Ral21]. EnCyCriS [ECPJ^{+}23, ESGJ^{+}23]. Engineering [AAY21a, AM23, BHC20, BHC21, BFS22, BBC^{+}20, BCL^{+}23b, BCL^{+}23a, CLG^{+}22, CS20, CSS20, DP23, ECPJ^{+}23, ESGJ^{+}23, FTP^{+}23, FBS22, For22, FM21, FU21, GMD24, HPSV22, HPSV23, HX20, JTNBC24, Kam23a, Kam23b, KFG^{+}21, KT23, KBT23, KPS20, LTG^{+}21, MZHS21, Med21, NSJ^{+}22, NSBC^{+}23, PD23, PK22, PPPC20, SCAB23, SABB24, SWG21, SKP20, Sol20a, Sol20b, Sol20c, Sol21a, Sol21b, TFTS20, VFHD20, VHFD21, WGW22, JJL23, SPC22, YAT^{+}21]. Ensemble [BBC^{+}20]. Ensemble-Based [BBC^{+}20]. Enterprise [Rog22a]. Equality [CSS20, JJL23]. Equity [FM23]. ESEC [KMS23]. ESEC/FSE [KMS23]. Essence [Leh23]. Ethical [ED24]. Evaluation [KMS23, VPE^{+}20]. Event [TRH^{+}20]. Evolution [Tzo20, Ven22]. Example [BHC20]. Experience [ED22b, TFTS20]. Experiences [Moh23]. Experimentation [Leh23]. Exploratory [Ner23]. Exploring [Leh23]. Export [Sch22b]. Exports [Sch22a, Sch22b, Sch23b]. EyeMovements [ABS23].

Fabric [Jah22]. Facing [ED22d]. Factors [Mba23]. Fellow [Pfa21b, Pfa21a, Pfa21c]. Fifth [FU21]. Finding [HKVD21, YWL^{+}21]. fine [PGG21]. fine-grained [PGG21]. First [ED23a, AAY21a, FTP^{+}23]. Formula [Sch23c]. Fostering [BCL^{+}23b]. Fourth [AM23, BCS^{+}23]. Framework [WKJ^{+}23, XdSBT24]. Freedom [Sch22a, Sch22b, Sch23b, Sch23a]. Frontends [TM22]. FSE [KMS23]. fun [BCL^{+}23a]. Functions [YAP23]. Fuzzing [GP23].

Gail [Sol23b]. Gain [ED23b]. Gains [Med21, SKP20, Sol20a, Sol20b, Sol20c, Sol21a, Sol21b]. Game [BCL^{+}23b]. Games [BCL^{+}23a]. GE [CSS20, JJL23]. Gender [CSS20, JJL23]. Generation [KT23, KBT23, PK22]. Generative [ED23c]. generic [RK21]. Genetic [Bro21, LWP^{+}20, LNP^{+}23]. Go [ED22a]. Golden [PPPC20]. good [PGM20]. Grail [ED22b]. grained [PGG21]. Great [ED22b]. Gregor [Rog22a]. Grundy [Sol23c]. Guidelines [WGB^{+}22].

HCSE&CS [CLG^{+}22]. Hohpe [Rog22a].

Holy [ED22b]. Home [KS23]. Human [CLG^{+}22, ED22c, Mba23]. Hybrid [Sch23c]. Hype [ED22a]. Hypergrowth [ED23a].

ICGSE [TRH^{+}20]. ICSA [LCA^{+}23]. ICSE [SABB24, Bro21, JJL23, LWP^{+}20, LNP^{+}23]. ICSSP [TRH^{+}20]. ICSSP-ICGSE [TRH^{+}20]. Identification [San23]. IEEE [AAC^{+}24, SABB24]. IEEE/ACM [AAC^{+}24]. Implementations [TM22]. Implementing [Rog22b]. Implications [Kam23b, KS23]. Improvement [Bro21, LWP^{+}20, LNP^{+}23]. Improving [Shi21]. Inclusion [FM23, JJL23]. Incremental [PK23]. Induced [Kam23a, KS23]. Industrial [MZHS21]. Industry [ED22a, ED22c, ED22d, ED23a, ED23b, ED23d, KS23, MBL21, WGB^{+}22]. Industry-Academia [MBL21]. Industry-Relevant [WGB^{+}22]. Influential [Wil20]. Infrastructure [CS20]. Initiative [RR20, Sch23a, TRK^{+}21]. inspiration [BCL^{+}23a]. Int [HPSV22, HPSV23, VFHD20, VHFD21]. Integration [Rog22a]. Intelligence [HPSV22, HPSV23, Sch20, VFHD20, VHFD21, YAT^{+}21]. InteNSE [JTNBC24]. Intensive [SGHJ20, WHW23]. interactions [Mau23]. International [AAY21a, ABS23, AAC^{+}24, AM23, BCS^{+}23, ECPJ^{+}23, ESGJ^{+}23, FU21, KFG^{+}21, RDF^{+}20, SABB24, Sch22a, Sch22b, Sch23b, Sch23a, SWG21, SGHJ20, WGW22, WHW23, XKD22, YAT^{+}21]. Internet [DMM21, NSBC^{+}23]. Interpretability [JTNBC24]. Intersections [GP23]. Interview [Pfa21b, Pfa21a, Pfa21c, Sol23a, Sol23b, Sol23c, Sol22a, Sol22b]. Introduction [Sch22a]. IoT4Emergency [DMM21]. Issue [San23].

Java [APA^{+}20, SNA^{+}23, WKJ^{+}23, YAP23, YWL^{+}21]. John [Sol23c]. Joint [TRH^{+}20]. Journal [Med21]. Journal-Centric [Med21]. JPF [BA23, ZA21, WKJ^{+}23]. jpf-logic [WKJ^{+}23]. Just [PK23].

Key [ED22b, ED22c, ED23d]. Kim [Sol22a].

Language [DP23, PD23]. Language-based [DP23, PD23]. Latin [SPC22]. Laurie [Pfa21a]. Layoff [ED23a]. Learned [FM21]. Learning [BCS^{+}23, Kam23a, Moh23]. Lessons [FM21]. libraries [RK21]. License [HKVD21]. Licensing [Sch22b]. Light [Leh23]. Light-Weight [Leh23]. Lionel [Pfa21b]. localization [RK21]. located [AAC^{+}24, KMS23]. logging [Mau23]. Logic [WKJ^{+}23, WKJ^{+}23]. LTD [XdSBT24].

made [Di 20]. Major [ED22d].

MALTESQUE

[FPA^{+}20, CXXS23, FCLA22].

Management

[DMM21, Kam23b, XdSBT24]. Managing [WCM^{+}23]. Mean [Sch20]. Measure [ED23d]. Measuring [ED23d]. Messaging [Rog22a]. Metamorphic [XKD22]. method [Mau23]. Metrics [ED23d]. Micro [TM22]. Micro-Frontends [TM22]. Microservices [Ham23]. Minimum [Leh23]. Miryung [Sol22a]. Model [AAY^{+}21b, PHD21]. Model-driven [AAY^{+}21b]. Modeling [LCA^{+}23]. Models [Tzo20, RK21]. Modern [BBC^{+}20]. Monolithic [Ham23]. motivation [BCL^{+}23a]. MSR4P&S [VFC23]. Multithreaded [PK23]. Murphy [Sol23b].

Nachiappan [Pfa21c]. Nagappan [Pfa21c]. Natural [DP23, PD23]. Needed [ED22b]. Networks [BHC21]. Neural [JTNBC24]. Neurodiversity [ED23b]. Next [KT23, KBT23, PK22]. NLBSE [DP23, PD23]. NLP [ARR^{+}22]. NLP-SEA [ARR^{+}22]. Not-So-Influential [Wil20].

Objects [Mau23]. One [ED22b]. Open [KFG^{+}21, San23, PGM20]. OpenAI [ED24]. Outbreak [BFS22]. Outstanding [Shi21].

Pains [Med21, SKP20, Sol20a, Sol20b, Sol20c, Sol21a, Sol21b]. Pairs [PK23]. Pandemic [Kam23a, KS23]. Pandemic-Induced [Kam23a, KS23]. Paper [RR20]. Part [Sch20, Sch22a, Sch22b, Sch23b, Sch23a]. Participant [Bro21]. Passages [Gro20a, Gro20b, Gro20c, Gro20d, Gro21a, Gro21b, Gro21c, Gro21d, Gro22a, Gro22b, Gro22d, Gro22c, Gro23a, Gro23b, Gro23c, Gro23d, Gro24]. Past [ED22a]. Pathfinder [APA^{+}20, SNA^{+}23, YWL^{+}21]. Patterns [ECN20, Rog22a]. Peer [HX20, Med21, Pez20, RR20, SKP20, Sol20a, Sol20b, Sol20c, Sol21a, Sol21b, PGM20]. Peer-Reviewing [SKP20, Sol20a, Sol20b, Sol20c, Sol21a, Sol21b]. Perceptions [PGM20]. Personal [Bro21]. Perspective [Med21]. Perspectives [HX20, Kam23b]. Physical [NSJ^{+}22, NSBC^{+}23]. Pitfalls [TM22]. Plans [Tzo20]. platform [Mau23]. Platforms [BBC^{+}20, SGHJ20]. Polluter [YWL^{+}21]. Practice [Gla20, KFG^{+}21, MZHS21]. Practices [Moh23]. Praise [Gla20]. prediction [PGG21]. Prejudice [Pez20]. Prerequisites [Leh23]. Principles [TM22]. Process [TRK^{+}21]. Product [dS23]. Products [ED22b]. Program [TRK^{+}21]. Programming [ABS23]. Programs [PK23]. Project [San23, BHC20]. Promoting [SCAB23]. Properties [WKJ^{+}23]. Public [Neu20a, Neu20b, Neu20c, Neu20d, Neu21a, Neu21b, Neu21c, Neu21d, Neu22a, Neu22b, Neu22c, Neu22d, Neu23a, Neu23b, Neu23c, Neu23d, Neu24].

Q [AAY21a, AM23, AAY^{+}21b]. Q-SE [AAY21a, AM23, AAY^{+}21b]. Qualities [SBT^{+}20]. Quality [NSJ^{+}22, NSBC^{+}23, RR20]. Quantum [AAY21a, AAY^{+}21b, AM23, FTP^{+}23, PPPC20]. Questions [KFG^{+}21].

Rapid [KFG^{+}21]. RCoSE [KFG^{+}21]. Readability [PHD21]. Realizing [YAT^{+}21]. Reflection [TRK^{+}21, Bro21]. Reflections [PHD21, Ven22]. Regression [Shi21]. Released [Ral21]. Relevant [WGB^{+}22]. Reliability [Shi21]. Report [AAY^{+}21b, AM23, CLG^{+}22, CS20, ECPJ^{+}23, ESGJ^{+}23, FTP^{+}23, GMD24, GHMS21, HPSV22, HPSV23, KFG^{+}21, KT23, KBT23, LCA^{+}23, NSJ^{+}22, NSBC^{+}23, PK22, RDF^{+}20, RR20, SABB24, SBT^{+}20, TFTS20, VFHD20, VHFD21, VPE^{+}20]. Reporting [SCAB23]. Reports [Niu20]. Requirements [GMD24]. Research [BHC20, BHC21, BCL^{+}23b, KFG^{+}21, MZHS21, MBL21, WGB^{+}22, WCM^{+}23]. Researching [Moh23]. Retrieval [RK21]. Retrospective [HKVD21, PM21, RK21, PGG21]. Review [ECN20, Pez20, RR20, Rog22a, Rog22b, PGM20]. Reviewing [HX20, Med21, SKP20, Sol20a, Sol20b, Sol20c, Sol21a, Sol21b]. Revolution [ED24]. Risk [Mba23]. Risks [Neu20a, Neu20b, Neu20c, Neu20d, Neu21a, Neu21b, Neu21c, Neu21d, Neu22a, Neu22b, Neu22c, Neu22d, Neu23a, Neu23b, Neu23c, Neu23d, Neu24]. Robustness [JTNBC24].

SBST [GP23]. Scalable [PK23]. School [SPC22]. Science [Ven22]. SCRUM [Ner23]. SE [AAY21a, AM23, AAY^{+}21b, PGM20]. SEA [ARR^{+}22, NSBC^{+}23]. SEA4DQ’21 [NSJ^{+}22]. Search [AAY^{+}21b]. Search-driven [AAY^{+}21b]. Second [TFTS20]. SEConfig [CS20]. Security [CLG^{+}22, Mba23, MGW20]. SEENG [KT23, KBT23, PK22]. Selection [KMS23, VPE^{+}20]. Self [WGB^{+}22, WCM^{+}23]. Self-Adaptation [WGB^{+}22]. Self-Adaptive [WCM^{+}23].

Seminar [FTP^{+}23]. SEmotion2020 [FU21]. Serious [BCL^{+}23b]. SERP4IoT’21 [GHMS21]. Service [Mau23]. Service-Action-Objects [Mau23]. SESoS [SABB24]. Shadow [TRK^{+}21]. Side [ED22c]. SIGSOFT [SGHJ20, Sol22a, Sol22b, Sol23a, Sol23b, Sol23c, Niu20, RR20, Ral21, Shi21, WHW23]. Simpler [PHD21]. Skill [San23]. Slow [ED22a]. Smallsats [Sch23b]. Smart [RDF^{+}20]. Social [dS23]. Socially [FM22]. Software [AAY21a, AAY^{+}21b, AAC^{+}24, AM23, BHC20, BHC21, BFS22, BBC^{+}20, BCL^{+}23b, BCL^{+}23a, CLG^{+}22, CS20, CSS20, DP23, ED22a, ED22b, ED22d, ED23b, ED23d, FTP^{+}23, FBS22, For22, FU21, Gla20, HKVD21, HPSV22, HPSV23, HX20, JTNBC24, JJL23, Kam23a, Kam23b, KS23, KFG^{+}21, KPS20, LCA^{+}23, LTG^{+}21, MZHS21, Med21, MGW20, Moh23, Ner23, NSJ^{+}22, NSBC^{+}23, PD23, PK22, PPPC20, PHD21, SCAB23, SABB24, SBT^{+}20, SWG21, SGHJ20, SKP20, Sol20a, Sol20b, Sol20c, Sol21a, Sol21b, SPC22, TFTS20, VFHD20, VHFD21, WGW22, WHW23, YAT^{+}21, dS23, Di 20, RK21, KT23, KBT23]. Software-Intensive [SGHJ20, WHW23]. Solutions [Rog22a]. Source [San23, PGG21]. Standards [Ral21]. Start [SGHJ20]. Start-ups [SGHJ20]. State [KFG^{+}21]. Status [RR20]. Strategies [FM22]. String [YAP23]. Students [Sch22a, Sch22b, Sch23b, Sch23a]. study [RK21]. STVR [HX20]. Success [ED22b, ED23d]. Suite [Shi21]. Summary [ABS23, ARR^{+}22, AAC^{+}24, BCS^{+}23, CPGP22, CXXS23, CSS20, DP23, FCLA22, FPA^{+}20, JTNBC24, JT22, PD23, SWG21, SPC22, TRH^{+}20, VFC23, WGW22, XKD22]. Support [WGB^{+}22, YAP23]. Supporting [San23]. Sustainability [dS23]. Sustainable [BHC20, WHW23, dS23]. Synergies [YAT^{+}21]. sYstem [AAC^{+}24]. Systems [ED23c, ECPJ^{+}23, ESGJ^{+}23, Ham23, NSJ^{+}22, NSBC^{+}23, SCAB23, San23, SABB24, WCM^{+}23, Mau23]. Systems-of-Systems [SCAB23, SABB24]. Systems/Internet [NSBC^{+}23].

Talent [ED22c]. Talk [Sch20]. Task [San23]. Task-driven [San23]. TC4JPF [ZA21]. Teaching [Kam23a]. Tech [ED22c, ED23a]. Technical [XdSBT24]. Tell [ED23c]. Temporal [WKJ^{+}23]. Tenth [ABS23]. Test [KMS23, Shi21, Tzo20, VPE^{+}20]. Test-Suite [Shi21]. Testing [BCS^{+}23, Ner23, Shi21, XKD22]. Tests [YWL^{+}21]. text [RK21]. Theia [BA23]. their [Kam23b, SBT^{+}20]. Things [DMM21, NSBC^{+}23]. Third [JJL23, WGW22]. Thoughts [Wil20]. Threads [PK23]. Topics [FBS22, For22]. TORACLE [JT22]. Trace [BA23, ZA21]. Traces [BA23, ZA21]. Tracking [San23]. Transferring [Moh23]. Transformations [Shi21]. Transforming [Ham23]. Travel [Niu20]. Trending [FBS22, For22]. Trust [Pez20]. try [Sch23c]. Tutorial [LCA^{+}23]. two [SBT^{+}20].

Uncertainty [WCM^{+}23]. Understanding [MGW20, WCM^{+}23]. University [FM22]. Unveiling [Mba23]. ups [SGHJ20]. Use [Ner23]. Used [ED23d]. Using [BA23, YWL^{+}21, ZA21, XdSBT24].

V [Rog22b]. verifiable [AAC^{+}24].

Verification [PK23]. Vernon [Rog22b]. via [Shi21]. Viewer [BA23]. Violations [HKVD21]. Visualize [BA23, ZA21]. vs [GP23].

Wave [ED23a, ED23c]. Weight [Leh23]. Wider [YAP23]. Williams [Pfa21a]. Within [Mba23]. Woolf [Rog22a]. Work [KS23]. Working [FTP^{+}23]. Workshop [AAY21a, ABS23, AAC^{+}24, AM23, APA^{+}20, BCS^{+}23, Bro21, CLG^{+}22, CS20, CSS20, DP23, ECPJ^{+}23, ESGJ^{+}23, FU21, GMD24, HPSV22, HPSV23, JJL23, KMS23, KFG^{+}21, PD23, RDF^{+}20, SABB24, SNA^{+}23, SWG21, SGHJ20, TFTS20, VFHD20, VHFD21, WGW22, WHW23, XKD22, YAT^{+}21, SBT^{+}20, AAY^{+}21b, ARR^{+}22, CPGP22, CXXS23, FCLA22, FPA^{+}20, GHMS21, JT22, KT23, KBT23, NSJ^{+}22, PK22, VFC23, VPE^{+}20]. Worldwide [KPS20]. would [PGM20]. WSQSE [FTP^{+}23].

Xia [Sol22b]. Xin [Sol22b].

Years [ED23a].

References

Arbab:2024:SIW

[AAC^{+}24] Farhad Arbab, Marco Autili, Federico Ciccozzi, Pascal Poizat, and Massimo Tivoli. Summary of the 5th International Workshop on Automated and verifiable Software sYstem DEvelopment (ASYDE)co-located with the 38th IEEE/ACM ASE 2023. ACM SIGSOFT Software Engineering Notes, 49(1): 24–26, January 2024. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3635439.3635444.

Abreu:2021:FIW

[AAY21a] Rui Abreu, Shaukat Ali, and Tao Yue. First International Workshop on Quantum Software Engineering (Q-SE 2020). ACM SIGSOFT Software Engineering Notes, 46 (2):30–32, April 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3448992.3449000.

Abreu:2021:QSM

[AAY^{+}21b] Rui Abreu, Shaukat Ali, Tao Yue, Michael Felderer, and Iaakov Exman. Quantum software: Model-driven or search-driven? A Q-SE 2021 Workshop report. ACM SIGSOFT Software Engineering Notes, 46(4): 23–25, October 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3485952.3485958.

AlMadi:2023:STI

[ABS23] Naser Al Madi, Teresa Busjahn, and Bonita Sharif. Summary of the Tenth International Workshop on EyeMovements in Programming (EMIP 2022). ACM SIGSOFT Software Engineering Notes, 48(1): 79–80, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573094.

Arcaini:2023:RFI

[AM23] Paolo Arcaini and Andriy Miranskyy. Report of the Fourth International Workshop on Quantum Software Engineering (Q-SE 2023). ACM SIGSOFT Software Engineering Notes, 48 (4):64–65, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617958.

Artho:2020:JPW

[APA^{+}20] Cyrille Artho, Quoc-Sang Phan, Peter Aldous, Alyas Almaawi, Lucas Bang, Lasse Berglund, Tevfik Bultan, Zhenbang Chen, Hayes Converse, Wei Dong, William Eiers, Milos Gligoric, Simon Goldsmith, Lars Grunske, Joshua Hooker, Ismet Burak Kadron, Timo Kehrer, Sarfraz Khurshid, Xuan-Bach D. Le, David Lo, Eric Mercer, Sasa Misailovic, Egor Namakonov, Hoang Lam Nguyen, Yannic Noller, Benjamin Ogles, Rohan Padhye, Pavel Parizek, Corina S. Pasareanu, S. Jacob Powell, Seemanta Saha, Koushik Sen, Elena Sherman, Kyle Storey, Minxing Tang, Willem Visser, Ji Wang, and Hengbiao Yu. The Java Pathfinder Workshop 2019. ACM SIGSOFT Software Engineering Notes, 45(2):20–22, April 2020.

URL https://dl.acm.org/doi/
10.1145/3385678.3385685.

Anwar:2022:NSW

[ARR^{+}22] Sajid Anwar, Abdul Rauf, Muhammad Ramzan, Imran Razzak, and Mehrdad Saadatmad. NLP-SEA 2021 Workshop summary. ACM SIGSOFT Software Engineering Notes, 47(2): 15–16, April 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3520273.3520279.

Besseling:2023:UTT

[BA23] Johan Besseling and Cyrille Artho. Using Theia trace viewer to visualize JPF traces. ACM SIGSOFT Software Engineering Notes, 48(1):22–26, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

                                        acm.org/doi/10.1145/3573074.
                                        3573081.

Brogi:2020:EBS

[BBC^{+}20] Antonio Brogi, Antonio Bucchiarone, Rafael Capilla, Pooyan Jamshidi, Maurizio Leotta, Zoltán Ádám Mann, Marina Mongiello, and Francesco Nocera. Ensemblebased software engineering for modern computing platforms. ACM SIGSOFT Software Engineering Notes, 45(1):28–30, January 2020. URL https://dl.

                                        acm.org/doi/10.1145/3375572.
                                        3375580.

Bucchiarone:2023:GSE

[BCL^{+}23a] Antonio Bucchiarone, Kendra M. L. Cooper, Dayi Lin, Edward F. Melcer, and Kelvin Sung. Games and software engineering: Engineering fun, inspiration, and motivation. ACM SIGSOFT Software Engineering Notes, 48 (1):85–89, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573096.

Bucchiarone:2023:FCA

[BCL^{+}23b] Antonio Bucchiarone, Kendra M. L. Cooper, Dayi Lin, Adam Smith, and Vanissa Wanick. Fostering collaboration and advancing research in software engineering and game development for serious contexts. ACM SIGSOFT Software Engineering Notes, 48 (4):46–50, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617955.

Biagiola:2023:SFI

[BCS^{+}23] Matteo Biagiola, Nicolás Cardozo, Donghwan Shin, Foutse Khomh, Andrea Stocco, and Vincenzo Riccio. Summary of the Fourth International Workshop on Deep Learning for Testing and Testing for Deep Learning (DeepTest 2023). ACM SIGSOFT Software Engineering Notes, 48(4): 39–40, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617953.

Breitenbucher:2022:SEA

[BFS22] Uwe Breitenbücher, Stefano Forti, and Jacopo Soldani. Software engineering after the COVID-19 outbreak. ACM SIGSOFT Software Engineering Notes, 47(4):7, October 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561846.3561848.

Bainomugisha:2020:SCB

[BHC20] Engineer Bainomugisha, Regina Hebig, and Michel Chaudron. Sustainable capacity building in software engineering research in Africa: The example of the BRIGHT Project. ACM SIGSOFT Software Engineering Notes, 45(3):18–20, July 2020.

URL https://dl.acm.org/doi/
10.1145/3402127.3402135.

Bainomugisha:2021:ESE

[BHC21] Engineer Bainomugisha, Regina Hebig, and Michel R. V. Chaudron. Emerging software engineering research networks in (East) Africa. ACM SIGSOFT Software Engineering Notes, 46 (2):18–22, April 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3448992.3448996.

Brownlee:2021:GII

[Bro21] Alexander E. I. Brownlee. Genetic improvement @ ICSE 2021: Personal reflection of a workshop participant. ACM SIGSOFT Software Engineering Notes, 46 (4):28–30, October 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3485952.3485960.

Chhetri:2022:RWH

[CLG^{+}22] Mohan Baruwal Chhetri, Xiao Liu, Marthie Grobler, Thuong Hoang, Karen Renaud, and Jennifer McIntosh. Report on the 2nd Workshop on Human Centric Software Engineering & Cyber Security (HCSE&CS 2021). ACM SIGSOFT Software Engineering Notes, 47(2): 12–14, April 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3520273.3520278.

Cooper:2022:AWS

[CPGP22] Kendra M. L. Cooper, Fabio Petrillo, Yann-Gaël Guéhéneuc, and Cristiano Politowski. ASE4Games 2021 Workshop summary. ACM SIGSOFT Software Engineering Notes, 47(2):10–11, April 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3520273.3520277.

Cito:2020:SEI

[CS20] Jürgen Cito and Mark Santolucito. Software engineering for infrastructure and configuration (SEConfig) — workshop report. ACM SIGSOFT Software Engineering Notes, 45 (2):23–24, April 2020. URL

https://dl.acm.org/doi/10.
1145/3385678.3385686.

Crnkovic:2020:SWG

[CSS20] Ivica Crnkovic, Karina Kohl Silveira, and Sara Sprenkle. Summary of the 2nd Workshop on Gender Equality in Software Engineering (GE 2019). ACM SIGSOFT Software Engineering Notes, 45(3):25–27, July 2020.

URL https://dl.acm.org/doi/
10.1145/3402127.3402138.

Cordy:2023:MWS

[CXXS23] Maxime Cordy, Xiaofei Xie, Bowen Xu, and Bibi Stamatia. MaLTeSQuE 2022 Workshop summary. ACM SIGSOFT Software Engineering Notes, 48 (1):95–96, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573099.

DiCosmo:2020:ABS

[Di 20] Roberto Di Cosmo. Announcing biblatex-software: software citation made easy. ACM SIGSOFT Software Engineering Notes, 45 (4):22–23, October 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3417564.3417570.

Dugdale:2021:IIT

[DMM21] Julie Dugdale, Mahyar T. Moghaddam, and Henry Muccini. IoT4Emergency: Internet of Things for emergency management. ACM SIGSOFT Software Engineering Notes, 46(1): 33–36, February 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3437479.3437489.

DiSorbo:2023:SNL

[DP23] Andrea Di Sorbo and Sebastiano Panichella. Summary of the 1st Natural Language-based Software Engineering Workshop (NLBSE 2022). ACM SIGSOFT Software Engineering Notes, 48(1): 101–104, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573101.

deSouza:2023:SSA

[dS23] Ana Carolina Moises de Souza. Social sustainability approaches for a sustainable software product. ACM SIGSOFT Software Engineering Notes, 48(1): 38–43, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573085.

Erl:2020:BRC

[ECN20] Thomas Erl, Robert Cope, and Amin Naserpour. Book review — cloud computing design patterns. ACM SIGSOFT Software Engineering Notes, 45 (2):27, April 2020. URL

https://dl.acm.org/doi/10.
1145/3385678.3385690.

Esnoul:2023:RIWa

[ECPJ^{+}23] Coralie Esnoul, Ricardo Colomo-Palacios, Eunkyoung Jee, Sabarathinam Chockalingam, John Eidar Simensen, and Doo-Hwan Bae. Report on the 3rd International Workshop on Engineering and Cybersecurity of Critical Systems (EnCyCriS — 2022). ACM SIGSOFT Software Engineering Notes, 48(1): 81–84, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573095.

El-Deeb:2022:AAW

[ED22a] Ahmed El-Deeb. AI adoption: Why the software industry is slow to go past the hype? ACM SIGSOFT Software Engineering Notes, 47(4): 16–17, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3561846.3561851.

El-Deeb:2022:HGS

[ED22b] Ahmed El-Deeb. The holy grail of software products success: Great customer experience and the key elements needed to create one. ACM SIGSOFT Software Engineering Notes, 47(2):8–9, April 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3520273.3520276.

El-Deeb:2022:HST

[ED22c] Ahmed El-Deeb. The human side of the tech industry: Key drivers behind the tech talent dilemma. ACM SIGSOFT Software Engineering Notes, 47(1): 10–11, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3502771.3502775.

El-Deeb:2022:MCC

[ED22d] Ahmed El-Deeb. Major challenges currently facing the software industry. ACM SIGSOFT Software Engineering Notes, 47 (3):14–15, July 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3539814.3539818.

El-Deeb:2023:FTL

[ED23a] Ahmed El-Deeb. The first tech layoff wave after years of hypergrowth: How this affects the industry? ACM SIGSOFT Software Engineering Notes, 48 (1):4–5, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573075.

El-Deeb:2023:NWC

[ED23b] Ahmed El-Deeb. Neurodiversity:

What can the software industry gain from it? ACM SIGSOFT Software Engineering Notes, 48 (4):19, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617950.

El-Deeb:2023:RWG

[ED23c] Ahmed El-Deeb. The recent wave of generative AI systems: What does this tell us about what AI can do now? ACM SIGSOFT Software Engineering Notes, 48(3):13, July 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3599975.3599979.

El-Deeb:2023:WSI

[ED23d] Ahmed El-Deeb. What the software industry is measuring?: Key metrics used by software companies to measure success. ACM SIGSOFT Software Engineering Notes, 48(2): 10–11, April 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3587062.3587065.

El-Deeb:2024:BOC

[ED24] Ahmed El-Deeb. Behind OpenAI CEO dismissal: an ethical dilemma and a new AI revolution. ACM SIGSOFT Software Engineering Notes, 49(1): 11–12, January 2024. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3635439.3635442.

Esnoul:2023:RIWb

[ESGJ^{+}23] Coralie Esnoul, Mary Sánchez-Gordón, Eunkyoung Jee, Ricardo Colomo-Palacios, Sabarathinam Chockalingam, John Eidar Simensen, and Doo-Hwan Bae. Report on the 4th International Workshop on Engineering and Cybersecurity of Critical Systems (EnCy-CriS — 2023). ACM SIGSOFT Software Engineering Notes, 48 (4):41–45, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3617946.3617954.

Forti:2022:TTSa

[FBS22] Stefano Forti, Uwe Breitenbücher, and Jacopo Soldani. Trending topics in software engineering. ACM SIGSOFT Software Engineering Notes, 47(3): 20–21, July 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3539814.3539820.

Feitosa:2022:MWS

[FCLA22] Daniel Feitosa, Gemma Catolino, Valentina Lenarduzzi, and Apostolos Ampatzoglou. MaL-TeSQuE 2021 Workshop summary. ACM SIGSOFT Software Engineering Notes, 47(1): 15–17, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3502771.3502777.

Fraser:2021:ECL

[FM21] Steven Fraser and Dennis Mancl.

Engineering for chaos: Lessons learned from COVID-19. ACM SIGSOFT Software Engineering Notes, 46(2):25–27, April 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448992.3448998.

Fraser:2022:SSD

[FM22] Steven Fraser and Dennis Mancl.

Strategies for “Socially Distant”: University–company collaborations. ACM SIGSOFT Software Engineering Notes, 47(1): 12–14, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3502771.3502776.

Fraser:2023:DDE

[FM23] Steven Fraser and Dennis Mancl.

Dimensions of diversity, equity, and inclusion. ACM SIGSOFT Software Engineering Notes, 48 (2):18–21, April 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3587062.3587068.

Forti:2022:TTSb

[For22] Stefano Forti. Trending topics in software engineering (1). ACM SIGSOFT Software Engineering Notes, 47(4): 6, October 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3561846.3561847.

Fontana:2020:MWS

[FPA^{+}20] Francesca Arcelli Fontana, Gilles Perrouin, Apostolos Ampatzoglou, Mathieu Archer, Bartosz Walter, Maxime Cordy, Fabio Palomba, and Xavier Devroey. MALTESQUE 2019 Workshop summary. ACM SIGSOFT Software Engineering Notes, 45(1): 34–35, January 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3375572.3375582.

Felderer:2023:SEC

[FTP^{+}23] Michael Felderer, Davide Taibi, Fabio Palomba, Michael Epping, Malte Lochau, and Benjamin Weder. Software engineering challenges for quantum computing: Report from the First Working Seminar on Quantum Software Engineering (WSQSE 22). ACM SIGSOFT Software Engineering Notes, 48 (2):29–32, April 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3587062.3587071.

Fucci:2021:FIW

[FU21] Davide Fucci and Hidetake Uwano. Fifth International Workshop on Emotion Awareness in Software Engineering (SEmotion2020). ACM SIGSOFT Software Engineering Notes, 46(1): 28–29, February 2021. URL

https://dl.acm.org/doi/10.
1145/3437479.3437487.

Gueheneuc:2021:SWR

[GHMS21] Yann-Gaël Guéhéneuc, Shah Rukh Humayoun, Rodrigo Morales, and Rubén Saborido. SERP4IoT’21 Workshop report. ACM SIG-SOFT Software Engineering Notes, 46(4):26–27, October 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

acm.org/doi/10.1145/3485952.
3485959.

Glass:2020:PSP

[Gla20] Robert L. Glass. In praise of software practice. ACM SIGSOFT Software Engineering Notes, 45(2):4–6, April 2020.

URL https://dl.acm.org/doi/
10.1145/3385678.3385680.

Gervasi:2024:RWE

[GMD24] Vincenzo Gervasi, Alessandro Marchetto, and Maya Daneva. Report of the 8th Workshop on Empirical Requirements Engineering (EmpiRE 2023). ACM SIGSOFT Software Engineering Notes, 49(1):27–29, January 2024. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

acm.org/doi/10.1145/3635439.
3635445.

Guizzo:2023:FVS

[GP23] Giovani Guizzo and Sebastiano Panichella. Fuzzing vs SBST: Intersections & differences. ACM SIGSOFT Software Engineering Notes, 48(1): 105–107, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573102.

Groce:2020:Pa

[Gro20a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 45(1):4–5, January 2020.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3375572.3375573.

Groce:2020:Pb

[Gro20b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 45(2):3–4, April 2020.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3385678.3385679.

Groce:2020:Pc

[Gro20c] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 45(3):4–5, July 2020.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3402127.3402129.

Groce:2020:Pd

[Gro20d] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 45(4):1–10, October 2020.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3417564.3417565.

Groce:2021:Pa

[Gro21a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 46(1):9–12, February 2021.

URL https://dl.acm.org/doi/
10.1145/3437479.3437481.

Groce:2021:Pb

[Gro21b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 46(2):10, April 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3448992.3448993.

Groce:2021:Pc

[Gro21c] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 46(3):7, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468745.

Groce:2021:Pd

[Gro21d] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 46(4):7, October 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485952.3485953.

Groce:2022:Pa

[Gro22a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 47(1):4, January 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3502771.3502772.

Groce:2022:Pb

[Gro22b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 47(2):3, April 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3520273.3520274.

Groce:2022:Pd

[Gro22c] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 47(4):8, October 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561846.3561849.

Groce:2022:Pc

[Gro22d] Alex S. Groce. Passages. ACM SIGSOFT Software Engineering Notes, 47(3):4, July 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3539814.3539815.

Groce:2023:Pa

[Gro23a] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 48(1):6, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3573074.3573076.

Groce:2023:Pb

[Gro23b] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 48(2):9, April 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3587062.3587064.

Groce:2023:Pc

[Gro23c] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 48(3):8, July 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3599975.3599977.

Groce:2023:Pd

[Gro23d] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 48(4):13, October 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617946.3617948.

Groce:2024:P

[Gro24] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 49(1):10, January 2024. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635439.3635441.

Hamza:2023:TMS

[Ham23] Muhammad Hamza. Transforming monolithic systems to a microservices architecture. ACM SIGSOFT Software Engineering Notes, 48(1):67–69, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

acm.org/doi/10.1145/3573074.
3573091.

Heinonen:2023:CCC

[Hei23] Ava Heinonen. Component comprehension in context. ACM SIGSOFT Software Engineering Notes, 48(1):44–48, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

acm.org/doi/10.1145/3573074.
3573086.

Hemel:2021:FSL

[HKVD21] Armijn Hemel, Karl Trygve Kalleberg, Rob Vermaas, and Eelco Dolstra. Finding software license violations through binary code clone detection — a retrospective. ACM SIGSOFT Software Engineering Notes, 46 (3):24–25, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468752.

Henry:2022:RIW

[HPSV22] Julie Henry, Upsorn Praphamontripong, Camelia Serban, and Andreea Vescan. Report from the 3rd Int. Workshop on Education through Advanced Software Engineering and Artificial Intelligence (EASEAI ’21). ACM SIGSOFT Software Engineering Notes, 47 (1):22–24, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3502771.350279.

Henry:2023:RIW

[HPSV23] Julie Henry, Upsorn Praphamontripong, Camelia Serban, and Andreea Vescan. Report from the 4th Int. Workshop on Education through Advanced Software Engineering and Artificial Intelligence (EASEAI ’22). ACM SIGSOFT Software Engineering Notes, 48 (2):22–23, April 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3587062.3587069.

Hierons:2020:PRS

[HX20] Robert M. Hierons and Tao Xie. Peer reviewing in software engineering: Perspectives from STVR Co-Editors-in-Chief. ACM SIGSOFT Software Engineering Notes, 45(4):18, October 2020.

URL https://dl.acm.org/doi/
10.1145/3417564.3417568.

Jahns:2022:DFD

[Jah22] Veit Jahns. Data fabric and datafication. ACM SIGSOFT Software Engineering Notes, 47 (4):30–31, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3561846.3561854.

Jaccheri:2023:GTW

[JJL23] Letizia Jaccheri, Lafifa Jamal, and Valentina Lenarduzzi. GE@ICSE: Third Workshop on Gender Equality, Diversity, and Inclusion in Software Engineering. ACM SIGSOFT Software Engineering Notes, 48(1): 90, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573097.

Jahangirova:2022:TWS

[JT22] Gunel Jahangirova and Valerio Terragni. TORACLE 2021 Workshop summary. ACM SIGSOFT Software Engineering Notes, 47 (1):25, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3502771.3502780.

Jabbarvand:2024:SIR

[JTNBC24] Reyhaneh Jabbarvand, Saeid Tizpaz-Niari, Earl T. Barr, and Satish Chandra. Summary of the 1st Interpretability and Robustness in Neural Software Engineering (InteNSE 2023). ACM SIGSOFT Software Engineering Notes, 49(1): 30–33, January 2024. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3635439.3635446.

Kamthan:2023:CPI

[Kam23a] Pankaj Kamthan. On COVID-19 pandemic-induced attitudinal changes in software engineering teaching and learning. ACM SIGSOFT Software Engineering Notes, 48(1): 13–14, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573078.

Kamthan:2023:DPM

[Kam23b] Pankaj Kamthan. On the different perspectives of management in software engineering education and their implications. ACM SIGSOFT Software Engineering Notes, 48(3): 14–15, July 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3599975.3599980.

Krusche:2023:SEEb

[KBT23] Stephan Krusche, Jonathan Bell, and Bastian Tenbergen. Software Engineering Education for the Next Generation: SEENG 2023 Workshop report. ACM SIGSOFT Software Engineering Notes, 48(4): 66–69, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617959.

Konersmann:2021:RCS

[KFG^{+}21] Marco Konersmann, Brian Fitzgerald, Michael Goedicke, Helena Holmström Olsson, Jan Bosch, and Stephan Krusche. Rapid continuous software engineering — state of the practice and open research questions: Report on the 6th International Workshop on Rapid Continuous Software Engineering (RCoSE 2020). ACM SIGSOFT Software Engineering Notes, 46 (1):25–27, February 2021. URL

https://dl.acm.org/doi/10.
1145/3437479.3437486.

Kiss:2023:WAT

[KMS23] Akos Kiss, Beatriz Marín, and Mehrdad Saadatmand. 13th Workshop on Automating Test Case Design, Selection and Evaluation (A-TEST 2022) colocated with ESEC/FSE conference. ACM SIGSOFT Software Engineering Notes, 48(1): 76–78, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573093.

Kuhrmann:2020:SEW

[KPS20] Marco Kuhrmann, Dietmar Pfahl, and Jacopo Soldani. Software engineering worldwide. ACM SIGSOFT Software Engineering Notes, 45 (3):17, July 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3402127.3402134.

Kamthan:2023:ICP

[KS23] Pankaj Kamthan and Nazlie Shahmir. On the implications of COVID-19 pandemic-induced “Work from home” in the software industry. ACM SIGSOFT Software Engineering Notes, 48 (2):12–13, April 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3587062.3587066.

Krusche:2023:SEEa

[KT23] Stephan Krusche and Bastian Tenbergen. Software Engineering Education for the Next Generation SEENG 2022 Workshop report. ACM SIGSOFT Software Engineering Notes, 48(1): 112–116, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573104.

Latifaj:2023:RBM

[LCA^{+}23] Malvina Latifaj, Federico Ciccozzi, Muhammad Waseem Anwar, Kousar Aslam, and Ivano Malavolta. Report on the blended modeling for software architectures tutorial at ICSA 2023. ACM SIGSOFT Software Engineering Notes, 48(3): 20–23, July 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3599975.3599983.

Lehtela:2023:EEC

[Leh23]  Bettina Lehtelä.  Explor-

ing the essence of continuous experimentation: Minimum prerequisites for light-weight CE. ACM SIGSOFT Software Engineering Notes, 48(1): 49–53, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573087.

Langdon:2023:GII

[LNP^{+}23] William B. Langdon, Vesna Nowack, Justyna Petke, Markus Wagner, Hyeonseok Lee, Erik M. Fredericks, Gabin An, and Aymeric Blot. Genetic improvement @ ICSE 2023. ACM SIGSOFT Software Engineering Notes, 48(4):51–59, October 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

                                        acm.org/doi/10.1145/3617946.
                                        3617956.

Licorish:2021:SEA

[LTG^{+}21] Sherlock A. Licorish, Christoph Treude, John Grundy, Kelly Blincoe, Stephen MacDonell, Chakkrit Tantithamthavorn, Li Li, and Jean-Guy Schneider. Software engineering in Australasia. ACM SIGSOFT Software Engineering Notes, 46(2): 16–17, April 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3448992.3448995.

Langdon:2020:GII

[LWP^{+}20] William B. Langdon, Westley Weimer, Justyna Petke, Erik Fredericks, Seongmin Lee, Emily Winter, Michail Basios, Myra B. Cohen, Aymeric Blot, Markus Wagner, Bobby R. Bruce, Shin Yoo, Simos Gerasimou, Oliver Krauss, Yu Huang, and Michael Gerten. Genetic improvement @ ICSE 2020. ACM SIGSOFT Software Engineering Notes, 45 (4):24–30, October 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3417564.3417575.

Maus:2023:SAO

[Mau23] Adam Maus. Service-action-objects: a cross-platform method for logging interactions in digital systems. ACM SIGSOFT Software Engineering Notes, 48 (2):14–17, April 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3587062.3587067.

Mbaka:2023:TUE

[Mba23] Winnie Mbaka. Towards unveiling effects of human factors within security risk assessment. ACM SIGSOFT Software Engineering Notes, 48(1): 70–75, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573092.

Marijan:2021:IAR

[MBL21] Dusica Marijan, Chetan Bansal, and Tamara Lopez. Industryacademia research collaborations during and after COVID-19. ACM SIGSOFT Software Engineering Notes, 46(4): 19–22, October 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3485952.3485957.

Medvidovic:2021:PGP

[Med21] Nenad Medvidović. Pains and gains of peer reviewing in software engineering: a journalcentric perspective. ACM SIGSOFT Software Engineering Notes, 46(3):15–16, July 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3468744.3468748.

Mirakhorli:2020:USS

[MGW20] Mehdi Mirakhorli, Matthias Galster, and Laurie Williams. Understanding software security from design to deployment. ACM SIGSOFT Software Engineering Notes, 45(2):25–26, April 2020.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3385678.3385687.

Mohammadi:2023:RTB

[Moh23] Mahyar Mohammadi. Researching and transferring blended learning experiences to software development practices. ACM SIGSOFT Software Engineering Notes, 48(1):63–66, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

                                        acm.org/doi/10.1145/3573074.
                                        3573090.

Marijan:2021:BSE

[MZHS21] Dusica Marijan, Thomas Zimmermann, MyungJoo Ham, and Bran Selic. Bridging software engineering research and industrial practice. ACM SIGSOFT Software Engineering Notes, 46 (1):30–32, February 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3437479.3437488.

Neri:2023:UES

[Ner23] Giulia R. Neri. The use of exploratory software testing in SCRUM. ACM SIGSOFT Software Engineering Notes, 48(1): 59–62, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573089.

Neumann:2020:RPa

[Neu20a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 45 (1):6–11, January 2020. URL

https://dl.acm.org/doi/10.
1145/3375572.3375574.

Neumann:2020:RPb

[Neu20b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 45 (2):7–12, April 2020. URL

https://dl.acm.org/doi/10.
1145/3385678.3385682.

Neumann:2020:RPc

[Neu20c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 45 (3):6–11, July 2020. URL

https://dl.acm.org/doi/10.
1145/3402127.3402130.

Neumann:2020:RPd

[Neu20d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 45 (4):11–16, October 2020. URL

https://dl.acm.org/doi/10.
1145/3417564.3417566.

Neumann:2021:RPa

[Neu21a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 46(1): 13–18, February 2021. URL

https://dl.acm.org/doi/10.
1145/3437479.3437482.

Neumann:2021:RPb

[Neu21b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 46 (2):11–15, April 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3448992.3448994.

Neumann:2021:RPc

[Neu21c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 46 (3):8–13, July 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468746.

Neumann:2021:RPd

[Neu21d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 46 (4):8–13, October 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3485952.3485954.

Neumann:2022:RPa

[Neu22a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 47 (1):5–7, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3502771.3502773.

Neumann:2022:RPb

[Neu22b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 47 (2):4–7, April 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3520273.3520275.

Neumann:2022:RPc

[Neu22c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 47 (3):5–10, July 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3539814.3539816.

Neumann:2022:RPd

[Neu22d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 47 (4):9–15, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3561846.3561850.

Neumann:2023:RPa

[Neu23a] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 48 (1):7–12, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573077.

Neumann:2023:RPb

[Neu23b] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 48 (2):4–8, April 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3587062.3587063.

Neumann:2023:RPc

[Neu23c] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 48 (3):4–7, July 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3599975.3599976.

Neumann:2023:RPd

[Neu23d] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 48 (4):8–12, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617947.

Neumann:2024:RP

[Neu24] Peter G. Neumann. Risks to the public. ACM SIGSOFT Software Engineering Notes, 49 (1):4–9, January 2024. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3635439.3635440.

Niu:2020:RSC

[Niu20] Nan Niu. Reports from SIG-SOFT CAPS travel awardees of ASE 2019. ACM SIGSOFT Software Engineering Notes, 45 (2):19, April 2020. URL

https://dl.acm.org/doi/10.
1145/3385678.3385684.

Nguyen:2023:SEA

[NSBC^{+}23] Phu H. Nguyen, Sagar Sen, Beatriz Bretones-Cassoli, Nicolas Jourdan, and Maria Chiara Magnanini. Software engineering and AI for data quality in cyber-physical systems/Internet of Things — SEA4DQ’22 report. ACM SIGSOFT Software Engineering Notes, 48(1): 108–111, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573103.

Nguyen:2022:SEA

[NSJ^{+}22] Phu H. Nguyen, Sagar Sen, Nicolas Jourdan, Beatriz Cassoli, Per Myrseth, Mikel Armendia, and Odd Myklebust. Software engineering and AI for data quality in cyber-physical systems — SEA4DQ’21 Workshop report. ACM SIGSOFT Software Engineering Notes, 47 (1):26–29, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3502771.3502781.

Panichella:2023:SNL

[PD23] Sebastiano Panichella and Andrea Di Sorbo. Summary of the 2nd Natural Language-based Software Engineering Workshop (NLBSE 2023). ACM SIGSOFT Software Engineering Notes, 48 (4):60–63, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617957.

Pezze:2020:PRT

[Pez20] Mauro Pezzè. Peer review: Trust and prejudice. ACM SIGSOFT Software Engineering Notes, 45 (4):19–21, October 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3417564.3417569.

Pfahl:2021:ILW

[Pfa21a] Dietmar Pfahl. An interview with Laurie Williams — ACM Fellow 2020. ACM SIGSOFT Software Engineering Notes, 46 (3):19–20, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3468744.3468750.

Pfahl:2021:ILB

[Pfa21b] Dietmar Pfahl. An interview with Lionel Briand — ACM Fellow 2020. ACM SIGSOFT Software Engineering Notes, 46 (2):23–24, April 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3448992.3448997.

Pfahl:2021:INN

[Pfa21c] Dietmar Pfahl. An interview with Nachiappan Nagappan — ACM Fellow 2020. ACM SIGSOFT Software Engineering Notes, 46 (4):14–15, October 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3485952.3485955.

Pinzger:2021:CFG

[PGG21] Martin Pinzger, Emanuel Giger, and Harald C. Gall. Comparing fine-grained source code changes and code churn for bug prediction — a retrospective. ACM SIGSOFT Software Engineering Notes, 46(3):

21–23, July 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468751.

Prechelt:2020:DBG

[PGM20] Lutz Prechelt, Daniel Graziotin, and Daniel Mendez. Double-blind is good but open would be better: Perceptions of peer review in the SE community. ACM SIGSOFT Software Engineering Notes, 45(3):16, July 2020.

URL https://dl.acm.org/doi/
10.1145/3402127.3402133.

Posnett:2021:RSM

[PHD21] Daryl Posnett, Abram Hindle, and Premkumar Devanbu. Reflections on: a simpler model of software readability. ACM SIGSOFT Software Engineering Notes, 46 (3):30–32, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468754.

Peraire:2022:SEE

[PK22] Cécile Péraire and Stephan Krusche. Software engineering education for the next generation SEENG 2021 Workshop report. ACM SIGSOFT Software Engineering Notes, 47(1): 18–21, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3502771.3502778.

Parizek:2023:CJP

[PK23] Pavel Parízek and Filip Kliber. Checking just pairs of threads for efficient and scalable incremental verification of multithreaded programs. ACM SIGSOFT Software Engineering Notes, 48(1): 27–31, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573082.

Pagano:2021:HDD

[PM21] Dennis Pagano and Walid Maalej. How do developers blog?: a retrospective. ACM SIGSOFT Software Engineering Notes, 46 (3):26–29, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3468744.3468753.

Piattini:2020:QCN

[PPPC20] Mario Piattini, Guido Peterssen, and Ricardo Pérez-Castillo. Quantum computing: a new software engineering Golden Age. ACM SIGSOFT Software Engineering Notes, 45 (3):12–14, July 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3402127.3402131.

Ralph:2021:ASE

[Ral21] Paul Ralph. ACM SIGSOFT empirical standards released. ACM SIGSOFT Software Engineering Notes, 46(1):19, February 2021.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3437479.3437483.

Raibulet:2020:RIW

[RDF^{+}20] Claudia Raibulet, Khalil Drira, MariaGrazia Fugini, Genaina Nunes Rodrigues, Patrizio Pelliccione, and Tomás Bures. Report of the 2nd International Workshop on Context-aware Autonomous and Smart Architectures. ACM SIGSOFT Software Engineering Notes, 45(1): 14–17, January 2020. URL

https://dl.acm.org/doi/10.
1145/3375572.3375576.

Rao:2021:RSL

[RK21] Shivani Rao and Avinash Kak.

Retrieval from software libraries for bug localization: a comparative study of generic and composite text models — a retrospective. ACM SIGSOFT Software Engineering Notes, 46 (3):33–36, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468755.

Rogers:2022:BRE

[Rog22a] David S. Rogers. Book review — Enterprise Integration Patterns: Designing, Building and Deploying Messaging Solutions (by Gregor Hohpe and Bobby Woolf). ACM SIGSOFT Software Engineering Notes, 47 (4):34, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3561846.3561856.

Rogers:2022:BRI

[Rog22b] David S. Rogers. Book review: Implementing Domain-Driven Design (by V. Vernon). ACM SIGSOFT Software Engineering Notes, 47(3):24, July 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3539814.3539822.

Ralph:2020:ASP

[RR20] Paul Ralph and Romain Robbes. The ACM SIGSOFT paper and peer review quality initiative: Status report. ACM SIGSOFT Software Engineering Notes, 45 (2):17–18, April 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3385678.3385681.

Santos:2024:RAI

[SABB24] Rodrigo Santos, Pablo Antonino, Antonia Bertolino, and Doo-Hwan Bae. Report on the 11th ACM/IEEE International Workshop on Software Engineering for Systems-of-Systems and Software Ecosystems — SESoS@ICSE 2023. ACM SIGSOFT Software Engineering Notes, 49(1): 34–37, January 2024. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3635439.3635447.

Santos:2023:STD

[San23] Fabio Santos. Supporting the task-driven skill identification in open source project issue tracking systems. ACM SIGSOFT Software Engineering Notes, 48 (1):54–58, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573088.

Sentilles:2020:SQT

[SBT^{+}20] Severine Sentilles, Barry Boehm, Catia Trubiani, Xavier Franch, and Anne Koziole. Software qualities and their dependencies report on two editions of the workshop. ACM SIGSOFT Software Engineering Notes, 45 (1):31–33, January 2020. URL

https://dl.acm.org/doi/10.
1145/3375572.3375581.

Santos:2023:RDP

[SCAB23] Rodrigo Santos, Eleni Constantinou, Pablo Antonino, and Jan Bosch. Reporting on a decade of promoting software engineering for systemsof-systems and software ecosystems. ACM SIGSOFT Software Engineering Notes, 48(2): 24–28, April 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3587062.3587070.

Schaefer:2020:WDW

[Sch20] Robert Schaefer. What do we mean when we talk about artificial intelligence?: (part 2). ACM SIGSOFT Software Engineering Notes, 45(2):13–16, April 2020.

URL https://dl.acm.org/doi/
10.1145/3385678.3385683.

Schaefer:2022:AFDa

[Sch22a] Robert Schaefer. Academic freedom, deemed exports, and international students, Part 1: Introduction. ACM SIGSOFT Software Engineering Notes, 47 (1):8–9, January 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3502771.3502774.

Schaefer:2022:AFDb

[Sch22b] Robert Schaefer. Academic freedom, deemed exports and international students, Part 2: Export licensing and compliance. ACM SIGSOFT Software Engineering Notes, 47(3): 16–19, July 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3539814.3539819.

Schaefer:2023:AFI

[Sch23a] Robert Schaefer. Academic freedom and international students, Part 4: The China initiative. ACM SIGSOFT Software Engineering Notes, 48(3): 9–12, July 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3599975.3599978.

Schaefer:2023:AFD

[Sch23b] Robert Schaefer. Academic freedom, deemed exports and international students, Part 3: Smallsats and cubesats. ACM SIGSOFT Software Engineering Notes, 48(1):15–18, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https://dl.

acm.org/doi/10.1145/3573074.
3573079.

Schaefer:2023:FHE

[Sch23c] Robert Schaefer. Formula hybrid + electric: Because it’s crazy to try. ACM SIGSOFT Software Engineering Notes, 48 (4):14–18, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617949.

Smolander:2020:ASI

[SGHJ20] Kari Smolander, Paul Grünbacher, Sami Hyrynsalmi, and Slinger Jansen. ACM SIGSOFT International Workshop on Software-Intensive Business: Start-ups, Platforms, and Ecosystems. ACM SIGSOFT Software Engineering Notes, 45(1):18–20, January

2020. URL https://dl.acm.
org/doi/10.1145/3375572.3375577.

Shi:2021:SOD

[Shi21] August Shi. SIGSOFT Outstanding Doctoral Dissertation Award: Improving regression testing efficiency and reliability via test-suite transformations. ACM SIGSOFT Software Engineering Notes, 46(3): 17–18, July 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468749.

Soldani:2020:PGPa

[SKP20] Jacopo Soldani, Marco Kuhrmann, and Dietmar Pfahl. Pains and gains of peer-reviewing in software engineering. ACM SIGSOFT Software Engineering Notes, 45(1):12–13, January

                                        2020. URL https://dl.acm.
                                        org/doi/10.1145/3375572.3375575.

Sherman:2023:JPW

[SNA^{+}23] Elena Sherman, Yannic Noller, Cyrille Artho, Franck van Breugel, Anto Nanah Ji, John Kellerman, Parssa Khazra, Filip Kliber, Gaurang Kudale, Pavel Parízek, Corina S. Pasareanu, Ron Pressler, Matt Walker, Hongru Wang, and Qiuchen Yan. The Java Pathfinder Workshop 2022. ACM SIGSOFT Software Engineering Notes, 48(1): 19–21, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573080.

Soldani:2020:PGPb

[Sol20a]  Jacopo Soldani.  Pains and

gains of peer-reviewing in software engineering. ACM SIG-SOFT Software Engineering Notes, 45(2):6, April 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3385678.3385688.

Soldani:2020:PGPc

[Sol20b] Jacopo Soldani. Pains and gains of peer-reviewing in software engineering (2). ACM SIGSOFT Software Engineering Notes, 45(3):15, July 2020.

URL https://dl.acm.org/doi/
10.1145/3402127.3402132.

Soldani:2020:PGPd

[Sol20c]  Jacopo Soldani.  Pains and

gains of peer-reviewing in software engineering (3). ACM SIGSOFT Software Engineering Notes, 45(4):17, October 2020.

URL https://dl.acm.org/doi/
10.1145/3417564.3417567.

Soldani:2021:PGPa

[Sol21a]  Jacopo Soldani.  Pains and

gains of peer-reviewing in software engineering (4). ACM SIGSOFT Software Engineering Notes, 46(1):8, February 2021.

URL https://dl.acm.org/doi/
10.1145/3437479.3437480.

Soldani:2021:PGPb

[Sol21b] Jacopo Soldani. Pains and gains of peer-reviewing in software engineering (5). ACM SIGSOFT Software Engineering Notes, 46(3):14, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3468744.3468747.

Soldani:2022:IMK

[Sol22a] Jacopo Soldani. An interview with Miryung Kim — 2022 SIG-SOFT Awardee. ACM SIGSOFT Software Engineering Notes, 47 (4):32–33, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3561846.3561855.

Soldani:2022:IXX

[Sol22b] Jacopo Soldani. An interview with Xin Xia — 2022 SIG-SOFT Awardee. ACM SIGSOFT Software Engineering Notes, 47 (3):22–23, July 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3539814.3539821.

Soldani:2023:ICC

[Sol23a]  Jacopo Soldani.  An inter-

view with Chunyang Chen — 2023 SIGSOFT Awardee. ACM SIGSOFT Software Engineering Notes, 48(3):16–17, July 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3599975.3599981.

Soldani:2023:IGM

[Sol23b] Jacopo Soldani. An interview with Gail Murphy — 2023 SIG-SOFT Awardee. ACM SIGSOFT Software Engineering Notes, 48 (4):37–38, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617952.

Soldani:2023:IJG

[Sol23c] Jacopo Soldani. An interview with John Grundy — 2023 SIG-SOFT Awardee. ACM SIGSOFT Software Engineering Notes, 48 (3):18–19, July 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3599975.3599982.

Steinmacher:2022:SBC

[SPC22] Igor Steinmacher Steinmacher, Rafael Prikladnicki, and Tayana Conte. Summary of the 12th Brazilian Conference on Software and the 1st Latin American School on Software Engineering. ACM SIGSOFT Software Engineering Notes, 47(3): 28–31, July 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3539814.3539824.

Shihab:2021:SIW

[SWG21] Emad Shihab, Stefan Wagner, and Marco Aurélio Gerosa. Summary of the 2nd International Workshop on Bots in Software Engineering (BotSE 2020). ACM SIGSOFT Software Engineering Notes, 46(1): 20–22, February 2021. URL

https://dl.acm.org/doi/10.
1145/3437479.3437484.

Tiwari:2020:ERS

[TFTS20] Saurabh Tiwari, Sheikh Umar Farooq, Ranjith Tharayil, and Paramvir Singh. An experience report on second workshop on emerging software engineering education. ACM SIGSOFT Software Engineering Notes, 45 (3):28–35, July 2020. URL

https://dl.acm.org/doi/10.
1145/3402127.3402139.

Taibi:2022:MFP

[TM22] Davide Taibi and Luca Mezzalira. Micro-frontends: Principles, implementations, and pitfalls. ACM SIGSOFT Software Engineering Notes, 47(4): 25–29, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3561846.3561853.

Tell:2020:SII

[TRH^{+}20] Paolo Tell, David Raffo, Liguo Huang, Igor Steinmacher, Ricardo Britto, Eray Tüzün, and Paul Clarke. Summary of the 1st ICSSP-ICGSE Joint Event. ACM SIGSOFT Software Engineering Notes, 45(4): 31–34, October 2020. URL

https://dl.acm.org/doi/10.
1145/3417564.3429719.

Thongtanunam:2021:SPC

[TRK^{+}21] Patanamon Thongtanunam, Ayushi Rastogi, Foutse Khomh, Serge Demeyer, Meiyappan Nagappan, Kelly Blincoe, and Gregorio Robles. Shadow program committee initiative: Process and reflection. ACM SIGSOFT Software Engineering Notes, 46(4): 16–18, October 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3485952.3485956.

Tzoref:2020:CEC

[Tzo20] Rachel Tzoref. Comprehension and evolution of combinatorial models and test plans. ACM SIGSOFT Software Engineering Notes, 45(3):23–24, July 2020.

URL https://dl.acm.org/doi/
10.1145/3402127.3402137.

Venkateswaran:2022:REC

[Ven22] Sreekrishnan Venkateswaran. Reflections on the evolution of computer science education. ACM SIGSOFT Software Engineering Notes, 47(3):11–13, July 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3539814.3539817.

Vidoni:2023:MWS

[VFC23] Melina Vidoni, Nicolás E. Díaz Ferreyra, and Zadia Codabux. MSR4P&S 2022 Workshop summary. ACM SIGSOFT Software Engineering Notes, 48(1): 97–100, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573100.

Vanderose:2020:RIW

[VFHD20] Benoît Vanderose, Benoît Frenay, Julie Henry, and Xavier Devroey. Report from the 1st Int. Workshop on Education through Advanced Software Engineering and Artificial Intelligence (EASEAI ’19). ACM SIGSOFT Software Engineering Notes, 45(1): 25–27, January 2020. URL

https://dl.acm.org/doi/10.
1145/3375572.3375579.

Vanderose:2021:RIW

[VHFD21] Benoît Vanderose, Julie Henry, Benoît Frénay, and Xavier Devroey. Report from the 2nd Int. Workshop on Education through Advanced Software Engineering and Artificial Intelligence (EASEAI ’20). ACM SIGSOFT Software Engineering Notes, 46 (2):28–29, April 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3448992.3448999.

Vos:2020:ATC

[VPE^{+}20] Tanja E. J. Vos, I. S. W. B. Prasetya, Sigrid Eldh, Sinem Getir, Ali Parsai, and Pekka Aho. Automating TEST case design, selection and evaluation report on 10 editions of A-TEST Workshop. ACM SIGSOFT Software Engineering Notes, 45(1): 21–24, January 2020. URL

https://dl.acm.org/doi/10.
1145/3375572.3375578.

Weyns:2023:TRA

[WCM^{+}23] Danny Weyns, Radu Calinescu, Raffaela Mirandola, Kenji Tei, Maribel Acosta, Amel Bennaceur, Nicolas Boltz, Tomas Bures, Javier Camara, Ada Diaconescu, Gregor Engels, Simos Gerasimou, Ilias Gerostathopoulos, Sinem Getir Yaman, Vincenzo Grassi, Sebastian Hahner, Emmanuel Letier, Marin Litoiu, Lina Marsso, Angelika Musil, Juergen Musil, Genaina Nunes Rodrigues, Diego Perez-Palacin, Federico Quin, Patrizia Scandurra, Antonio Vallecillo, and Andrea Zisman. Towards a research agenda for understanding and managing uncertainty in self-adaptive systems. ACM SIGSOFT Software Engineering Notes, 48(4): 20–36, October 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3617946.3617951.

Weyns:2022:GAS

[WGB^{+}22] Danny Weyns, Ilias Gerostathopoulos, Barbora Buhnova, Nicolás Cardozo, Emilia Cioroaica, Ivana Dusparic, Lars Grunske, Pooyan Jamshidi, Christine Julien, Judith Michael, Gabriel Moreno, Shiva Nejati, Patrizio Pelliccione, Federico Quin, Genaina Rodrigues, Bradley Schmerl, Marco Vieira, Thomas Vogel, and Rebekka Wohlrab. Guidelines for artifacts to support industry-relevant research on selfadaptation. ACM SIGSOFT Software Engineering Notes, 47 (4):18–24, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3561846.3561852.

Wagner:2022:STI

[WGW22] Stefan Wagner, Marco A. Gerosa, and Mairieli Wessel. Summary of the Third International Workshop on Bots in Software Engineering (BotSE 2021). ACM SIGSOFT Software Engineering Notes, 47(3): 25–27, July 2022. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3539814.3539823.

Werder:2023:AST

[WHW23] Karl Werder, Sami Hyrynsalmi, and Xiaofeng Wang. ACM SIG-SOFT towards sustainable software business: 5th International Workshop on Software-Intensive Business. ACM SIGSOFT Software Engineering Notes, 48(1): 91–94, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573098.

Wilson:2020:TIE

[Wil20] Greg Wilson. Thoughts from a not-so-influential educator. ACM SIGSOFT Software Engineering Notes, 45(3):21–22, July 2020.

                                        URL https://dl.acm.org/doi/
                                        10.1145/3402127.3402136.

Walker:2023:JLF

[WKJ^{+}23] Matt Walker, Parssa Khazra, Anto Nanah Ji, Hongru Wang, and Franck van Breugel. jpflogic: a framework for checking temporal logic properties of Java code. ACM SIGSOFT Software Engineering Notes, 48 (1):32–36, January 2023. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3573074.3573083.

Xavier:2024:ATD

[XdSBT24] Laerte Xavier, Rodrigo dos Santos, Sandalo Bessa, and Marco Tulio Valente. Agile technical debt management using the LTD framework. ACM SIGSOFT Software Engineering Notes, 49 (1):13–23, January 2024. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3635439.3635443.

Xie:2022:WSI

[XKD22] Xiaoyuan Xie, Upulee Kanewala, and Alastair F. Donaldson. Workshop summary: 7th International Workshop on Metamorphic Testing (MET 2022). ACM SIGSOFT Software Engineering Notes, 47 (4):35–36, October 2022. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3561846.3561857.

Yan:2023:TWS

[YAP23] Qiuchen Yan, Cyrille Artho, and Pavel Parízek. Towards wider support for Java string functions. ACM SIGSOFT Software Engineering Notes, 48(1): 37, January 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

URL https://dl.acm.org/doi/
10.1145/3573074.3573084.

Yoo:2021:IWR

[YAT^{+}21] Shin Yoo, Aldeida Aleti, Burak Turhan, Leandro L. Minku, Andriy Miranskyy, and Çetin Meriçli. The 8th International Workshop on Realizing Artificial Intelligence Synergies in Software Engineering. ACM SIGSOFT Software Engineering Notes, 46 (1):23–24, February 2021. URL

https://dl.acm.org/doi/10.
1145/3437479.3437485.

Yi:2021:FPT

[YWL^{+}21] Pu Yi, Anjiang Wei, Wing Lam, Tao Xie, and Darko Marinov. Finding polluter tests using Java PathFinder. ACM SIGSOFT Software Engineering Notes, 46 (3):37–41, July 2021. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3468744.3468756.

Zhou:2021:TUT

[ZA21] Yang Zhou and Cyrille Artho. TC4JPF: Using Trace Compass to visualize JPF traces. ACM SIGSOFT Software Engineering Notes, 46(3):42–46, July 2021. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3468744.3468757.